If the distance between (-2, -3) and (6, y) is 10, what is y?
Answer by Edwin McCravy(20054)

d = √(x2-x1)²+(y2-y1


d = √(6-(-2))²+(y-(-3))² = 10

√(6+2)²+(y+3)² = 10

√8²+(y+3)² = 10

√64+y²+6y+9 = 10

√73+y²+6y = 10

Square both sides:

(√73+y²+6y)² = (10)²

 73+y²+6y = 100

 y²+6y+73 = 100

 y²+6y-27 = 0

(y+9)(y-3) = 0

y+9=0;  y-3=0
  y=-9;   y=3

Edwin